Subject: Diffusion Morning Report 11/05/12
From: Ted Berg <tberg@stanford.edu>
Date: Mon, 05 Nov 2012 06:32:28 -0800
Mon, 05 Nov 2012 06:32:28 -0800
Good Morning All,
*1. New issues:User Jasonlin reported pressure control errors on Epi, 
but many users followed with none
* *2. Ongoing issues:*Tylan Poly rewriting recipes,** Tylannitride  
tested again waiting further instructions after alarms,  working on 
Badger interlocks,Epi regulator  assy installed but still 
creeps,*Thermco Nitride slow pumping speed.,**particles on back of Epi 
wafers*,*O2 in Silane? on Epi*,*Tylan 2 low growth,*
*3.  Resolved *:*Tylan 3 ok after repair**

*
					
	Date: 	11/5/12 	THICKNESS-Change at 		Priority
Tube Name/Number 	Status 				
Tylan 1 	*up* 	** 	** 	** 	**
Tylan 2 	*yellow* 	*still low dep* 	** 	** 	*2*
Tylan 3 	*up* 	** 	** 	** 	**
Tylan 4 	*up* 	** 	** 	** 	**
Tylan 5-BBr3 	*up* 	** 	** 	** 	**
Tylan 6-POCl3 	*up* 	** 	** 	** 	**
Tylan 7-TEOS2 	*up* 	** 	*14.39/30* 	** 	**
Tylan 8-SiGe 	*down* 	*waiting for users* 	*3.4/30* 	** 	**
Tylan 9-Poly 	*down* 	*rewriting recipes-testing revised recipes* 
*23.7/30* 	** 	**
Tylan 10-Nitride 	*down* 	*tested after sih4 alarm,testing more* 
*28.5/30* 	** 	**
Tylan 11-LTO 	*up* 	** 	*4.3/30* 	** 	**
Tylan 12-FGA 	*up* 	** 	** 	** 	**
Thermco 1 	*up* 	** 	** 	** 	**
Thermco 2 	*up* 	** 	** 	** 	**
Thermco 3 	*up* 	** 	** 	** 	**
Thermco4 	*up* 	** 	** 	** 	**
Thermco Poly1 	*up* 	** 	** 	*
* 	**
Thermco Poly2 	*up* 	** 	** 	** 	**
Thermco Nitride1 	*yellow* 	*slow pumping speed-poor uniformity-pump 
ordered* 	** 	** 	**
Thermco bpsg 	*up* 	** 	** 	** 	**
AMAT Epi Cham. A 	*yellow* 	*O2 in sih4?possible particles, H2 reg 
creep* 	** 	** 	*1*
AMAT Epi Cham. B 	*down* 	***At the county* 	** 	** 	**
TEL Spa 	*up* 	** 	** 	** 	**
Burn Box 	*up* 	** 	** 	** 	**
SEM 4160 	*up* 	** 	** 	** 	**
FGA2 	*up* 	** 	** 	** 	**
SiH4 for furnaces 	*on* 	*crossover ok* 	***changeover working 
well*** 	** 	**



Good Morning All,
    1. New issues:User Jasonlin reported pressure control errors on Epi, but many users followed with none
    2. Ongoing issues:Tylan Poly rewriting recipes,  Tylannitride  tested again waiting further instructions after alarms,  working on Badger interlocks,Epi regulator  assy installed but still creeps,Thermco Nitride slow pumping speed.,particles on back of Epi wafers,O2 in Silane? on Epi,Tylan 2 low growth,
    3.  Resolved :Tylan 3 ok after repair


                   
  Date: 11/5/12 THICKNESS-Change at    Priority
Tube Name/Number Status        
Tylan 1 up        
Tylan 2 yellow still low dep     2
Tylan 3 up        
Tylan 4 up        
Tylan 5-BBr3 up        
Tylan 6-POCl3 up        
Tylan 7-TEOS2 up   14.39/30    
Tylan 8-SiGe down waiting for users 3.4/30    
Tylan 9-Poly down rewriting recipes-testing revised recipes 23.7/30    
Tylan 10-Nitride down tested after sih4 alarm,testing more 28.5/30    
Tylan 11-LTO up   4.3/30    
Tylan 12-FGA up        
Thermco 1 up        
Thermco 2 up        
Thermco 3 up        
Thermco4 up        
Thermco Poly1 up    
 
Thermco Poly2 up        
Thermco Nitride1 yellow slow pumping speed-poor uniformity-pump ordered      
Thermco bpsg up        
AMAT Epi Cham. A yellow O2 in sih4?possible particles, H2 reg creep     1
AMAT Epi Cham. B down  At the county      
TEL Spa up        
Burn Box up         
SEM 4160 up        
FGA2 up        
SiH4 for furnaces on crossover ok  changeover working well